Advanced Accounting Ch 1 Q 2

Paxton, Inc., acquired 10 percent of Bethlehem Corporation on January 1, Year 10, for $190,000 and appropriately accounted for the investment using the fair-value method.  On January 1, Year 11, Paxton purchased an additional 30 percent of Bethlehem for $600,000 which resulted in significant influence over Bethlehem.  On that date, the fair value of Bethlehem’s common stock was $2,000,000 in total.  Bethlehem’s January 1, Year 11, book value equaled $1,850,000, although land was undervalued by $120,000.  Any additional excess fair value over Bethlehem’s book value was attributable to a trademark with an eight-year remaining life.  During Year 11, Bethlehem reported income of $300,000 and declared and paid dividends of $110,000.  Prepare the Year 11 journal entries for Paxton related to its investment in Bethlehem.  
Answer #1

Step 1: Find out how much the investor paid vs. how much should pay for the investment.

Should pay = 2,000,000 Fair value x 10% = 200,000

Actually paid = 190,000

200,000 - 190,000 = 10,000

Since the investor paid less than they actually should have paid, there will be a fair value adjustment of 10,000.


Step 2: Investor purchased 30% more.  2,000,000 x 30% = 600,000


Step 3: Total investment will be 600,000 second purchse + 200,000 inital purchase = 800,000


Step 4: Total book value = 1,850,000 book value year 11 + 120,000 undervalued land = 1,970,000


Step 5: 1,970,000 total book value x 40% investor's portion = 788,000 


Step 6: Find trademark amortization per year

800,000 - 788,000 = 12,000

12,000 / 8 years = 1,500 amortization per year


Step 7: Find investor's share of Net income and dividends

Investor's Net Income portion:   40% x 300,000 = 120,000

Investor's Dividend portion:  40% x 110,000 = 44,000


Step 8: Make a T Chart

__________________________________40%

200,000                        |

Initial purchase             | 

                                     |

600,000                        |       44,000

Additional purchase      |       Dividends

                                      |

120,000                         |       1,500

Net income                    |       amortization


Step 9: Prepare journal Entries


To record initial investment purchase with the fair value adjustment: 

Jan 1, Year 10Investment in Beth Corp200,000

         Invest. in Equity               Securities             200,000

         F.V. Adjustment             10,000


To record the additional portion purchased:

Jan 1, Year 11Investment in Beth Corp600,000

        Cash            600,000


To record the investor's portion of net income of $120,000, the dividends received of $44,000, and the trademark's amortization of $1,500:



Dec 31, Year 11Investment in Beth Corp
         Equity in Earnings           Affiliate
120,000
              120,000

                 

Cash
         Investment in Beth
44,000
               44,000




Equity in Earnings Aff.1,500

      Invest in Beth Corp               1,500












answered by: anonymous